A recessive allele causes a condition which is lethal before birth for homozygotes. The frequency of heterozygotes in the adult population is 1 in 100.

A. Are heterozygotes aware of their recessive allele?

B. What’s the chance that two healthy adults are both carriers for the condition?
Answer #1

ANS)

According to data Frequency of heterozygote = 1/100

= 0.01

A) No heterozygotes are not mindful of their recessive alleles. Individuals with genotype Aa (having both recessive and dominant alleles at the A locus) are heterozygotes. This is because heterozygotes don't exhibit symptoms of this disease. Because of lack of any symptoms heterozygotes (carriers) are often unaware of their recessive allele.

B)

Chance two healthy adults are both carriers of This disease can be determined by multiplicative rule (since both events are independent): P(A and B) = P(A)*P(B)


-->Probability two healthy adults are both carriers of the disease= 0.01*0.01 = 0.0001